# A Comprehensive Evaluation of Augmentations for Robust OOD Self-Supervised Contrastive Phonocardiogram Representation Learning

### 18 Mar 2024 v2

@@ Abstract
Despite the recent increase in research activity, deep-learning models have not yet been widely accepted in several real-world settings, such as medicine. The shortage of high-quality annotated data often hinders the development of robust and generalizable models, which do not suffer from degraded effectiveness when presented with newly-collected, out-of-distribution (OOD) datasets. Contrastive Self-Supervised Learning (SSL) offers a potential solution to labeled data scarcity, as it takes advantage of unlabeled data to increase model effectiveness and robustness. In this research, we propose applying contrastive SSL for detecting abnormalities in 1D phonocardiogram (PCG) samples by learning a generalized representation of the signal. Specifically, we perform an extensive comparative evaluation of a wide range of audio-based augmentations, evaluate trained classifiers on multiple datasets across different downstream tasks, and finally report on the impact of each augmentation in model training. We experimentally demonstrate that, depending on its training distribution, the effectiveness of a fully-supervised model can degrade up to 32% when evaluated on unseen data, while SSL models only lose up to 10% or even improve in some cases. We argue and experimentally demonstrate that, contrastive SSL pretraining can assist in providing robust classifiers which can generalize to unseen, OOD data, without relying on time- and labor-intensive annotation processes by medical experts. Furthermore, the proposed extensive evaluation protocol sheds light on the most promising and appropriate augmentations for robust PCG signal processing, by calculating their effect size on model training. Finally, we provide researchers and practitioners with a roadmap towards producing robust models for PCG classification, in addition to an open-source codebase for developing novel approaches.
{IEEEkeywords}
@@ 1 Introduction
[…]
Abnormal phonocardiogram (PCG) signals are often an important indicator of heart diseases or deficiencies linked to increased mortality [1]. The detection of such indicators in a non-invasive, accurate, and robust manner could assist towards prevention or timely treatment of such diseases, especially when access to healthcare services is limited.
Automated methods for detecting abnormal samples commonly rely on audio recordings; typically, audio is processed and then used to train classifiers that detect abnormal (i.e., non-healthy) samples and sometimes discriminate between different diseases (e.g., [2]). However, the audio quality of collected signals can be affected by several circumstances. While typical audio-related factors include microphone type or sampling rate, among others, for the case of PCG recordings additional factors can significantly affect the quality of their recording, such as placement of the microphone [3, 4], body size, skin, noise [5, 6], etc. The above cases can prove even more challenging, such as when screening an unborn child/fetus, where PCG can only be captured through the mother’s belly [7]. In addition to the quality of the recording conditions, the training of accurate ML models is heavily dependent on annotated data, the collection of which proves highly time- and cost-consuming. The above is especially important in the biomedical domain, where available samples are scarce to begin with.
As a result, classification models that have been trained on data originating from a single source or on low quality signals fail to generalize to previously unseen data distributions, which do not adhere to the i.i.d. assumption (i.e they are not independently and identically distributed), leading to degraded effectiveness [8, 9]. In this work, we propose training robust classifiers for PCG signal classification by employing contrastive self-supervised learning (SSL) and investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets. However, as the effectiveness of an backbone encoder trained via SSL heavily depends on the adopted augmentation policy [10], we also research the appropriate transformation to be applied on PCG signals. Specifically, we attempt to provide insight into the following research questions:
@@
  • If so, which augmentations or transformations lead to such representations, proving the most effective, and which actually harm training?
@@ 2.1 Deep Learning for Phonocardiogram Classification
The release of publicly available PCG datasets under the scope of the 2016 [11] and 2022 [12] PhysioNet [13] challenges, has a sparked a line of research in deep learning for PCG classification. The majority of previous works have mainly proposed either 2D or 1D CNN architectures for tackling this particular problem. The main difference lies in whether the proposed models are trained directly on the one dimensional recorded auscultation waves or on spectral images extracted from the audio recordings. For example, multiple recent papers [14, 15, 16, 17] propose combining mel-spectrogram images extracted from PCG signals with encoded patient demographic data, in a dual-branch CNN model. The authors find that the combination of the above representations leads to improved accuracy in the heart murmur detection task. Similarly, in another work utilizing frequency domain features, the authors of [18] design a hierarchical deep CNN architecture that extracts and combines features from multiple-scales of PCG Mel-Spectrograms. A more recent work, [19], proposes training an ensemble of 15 ResNets [20] with channel-wise attention mechanisms on similar spectrograms, for grading murmur intensity in PCG recordings.
Although few, 1D models also prove promising in the heart-murmur detection domain. In [21] the authors employ a transformer-based neural network and train it on 1D wavelet power features extracted from raw PCG signals. On the other hand, [22] proposes using a U-Net to predict heart states from the raw PCG signal and combine the predictions with a ResNet model for the final classification. In a different approach, the authors of [23] suggest that the combination of a 1D CNN-Bi-LSTM model coupled with attention mechanisms, trained in conjunction with a feature-based classifier leads to robust abnormal PCG signal detection.
In this work, we build upon our preliminary research [24] for the 2022 PhysioNet challenge. In particular, we demonstrate the value of adopting self-supervised contrastive learning for robust PCG classification and also report the top augmentations which should be implemented in such a framework. To our knowledge, this is the first study extensively researching the benefits of each augmentation, and their combinations, for contrastive SSL in 1D PCG signal processing. The aim of this study is to provide future researchers with best-practices regarding PCG signal augmentation, while also propose a structured protocol for the evaluation of developed methods.
@@ 2.2 Self-Supervised Learning in 1D biosignal classification
Self-supervised representation learning (SSL) [25] was initially proposed for diminishing the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks [26] to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding [27], the goal of Self-Supervised Contrastive Learning [28, 29] is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions [30]. The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which can take advantage of the small amount of data but also yield adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [31]. The authors of [32] propose augmenting multi-channel EEG signals by recombining several channels into additional channels which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying pre-defined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of [33] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in [34] where contrastive predictive coding is implemented on similar SSL-learned features to both predict sleep-stages and detect pathologies in multi-channel EEG signals. Finally, [35] proposed BENDR, which combines attention mechanisms [36] and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well [37], most of which focus on either emotion recognition or pathology detection. For example, in [38] the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the art results. In another paper studying emotion recognition from ECG signals, [39] introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [40], implementing attention mechanisms [41] or combining wavelet transformations and random crops of the signal [42] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, [43] propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models [44]. For example, [45] introduces NeuroPose, a system for 3D hand pose tracking which aims to alleviate the problems imposed due to different sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors [46] employ a contrastive loss to pretrain a model on EMG signals for hang-gesture recognition.
@@ 2.3 Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning [47]. To test and improve the generalizabilty prowess and robustness of deep learning models, several fields have emerged, such as Domain Adaptation [48], Transfer Learning [49], and Domain Generalization [50, 9, 51], which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data generating processes. Here, we reference prior works which fall under the above research fields.
@@
Domain adaptation (DA) methods are the most popular choice when aiming to produce models with improved generalization capabilities in 1D biosignal classification. Similar to Transfer Learning (TL), DA algorithms leverage the representation learning capabilities of models pretrained on diverse source data distributions and fine-tune them on previously unseen but similar target distributions. The end-result is a model which remains robust under the distribution shift between source and target data. In healthcare, DA methods are proposed for a plethora of biosignal classification tasks. In [52], the authors propose pre-training vanilla CNN networks on a large corpus of raw ECG signals and demonstrate their effectiveness when fine-tuned on a smaller dataset for arrhythmia detection. In another more recent work, He et al. [53] propose tackling the distribution shift between different arrhythmia detection ECG datasets, by employing three distribution alignment mechanisms. To be more specific, they first extract spatio-temporal features from preprocessed signals, pass them through a graph convolutional neural network and finally align the resulting representations between source and target domains. Many studies have researched the advantages of DA and TL in EEG classification as well [54]. To name a few, the authors of [55, 56] and [57] propose aligning source and target distributions via adversarial learning for motor image EEG classification and epilepsy detection respectively, while [58] investigates several state-of-the-art DA algorithms for EEG-based emotion recognition.
@@
Finally, Domain Generalization algorithms aim to tackle the model generalization at its core, as the main goal is to provide generalized models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of [59] propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, [60] and [61] propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of [9] introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from [62], for 12-lead ECG and 64-channel EEG classification.
@@ 3 Materials and Methods
For the self-supervised training of all models, we adapt the SimCLR [63] methodology for 1D PCG representation learning and then evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
@@ 3.1 Datasets & Downstream tasks
The PCG signals used in this study originate from a total of 5 distinct public datasets and are available as “wav” files. These include the Ephnogram [64], FPCGDB [65], Pascal [66], PhysioNet2016 [11], and PhysioNet2022 [12] databases. Here we describe each dataset separately and provide all necesary details regarding the preprocessing steps followed for the homogenization of all signals into a common format.
@@
The Ephnogram [64] database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals for noise research purposes. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled samples of the PCG signals for the SSL pretraining of our models. All PCG recordings were sampled at 8 kHz.
@@
The FPCGDB [65] or Fetal PCG Database, as its name suggests, is a dataset containing 26 20-minute fetal phonocardiographic signals, gathered from healthy women during their the remaining months of their pregnancy and sampled at 333 Hz. Similarly to the Ephnogram dataset, the collected signals were obtained for the development of fetal PCG signal simulation and fetal heart rate extraction algorithms. Hence, this dataset also does not contain any labeled samples for classification purposes. Furthermore, most recordings are somewhat corrupted by noise, either due to the pregnant human body which contains various noise sources, such as muscular movements or placental blood turbulence, or even due to the surrounding environment.
@@
The Pascal [66] Heart Sound Classification challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in both clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole [^2] Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced in the context of public challenges for PCG classification. Similarly to the Pascal challenge, the objective of the PhysioNet2016 [11] challenge aims at urging researchers to develop robust algorithms for differentiating normal and abnormal heart sound recordings but instead introduces a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 challenge [12] was designed for the development of algorithms which could detect heart murmur in PCG signals. Succedent to the 2016 challenge, the 2022 challenge contains multiple heart sound recordings, collected from multiple auscultation locations. The available data includes 3,163 recordings from 942 patients in total. Each patient has one or more PCG signals, recorded from one of the four heart valves or from a separate prominent auscultation location. Each of the recordings can be classified as ”present”, ”absent” or ”unknown”, meaning that heart murmur presence can be either detected, not detected or can not be determined (e.g due to noise or corrupted audio) respectively, rendering the challenge a multi-class classification problem. The sampling rate of the above signals is 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process as in [24] for converting all signals into a similar format. As a first step, to avoid transient noise artifacts, we initially discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether it derives from a normal heart sound recording or not [^3] In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage, we assign pseudolabels to each window, as described in Section 3.2. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can be then evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section 3.4.

### 05 Apr 2024 v3

# Which Augmentation Should I Use? An Empirical Investigation of Augmentations for Self-Supervised Phonocardiogram Representation Learning

@@ Abstract
Despite the recent increase in research activity, deep-learning models have not yet been widely accepted in several real-world settings, such as medicine. The shortage of high-quality annotated data often hinders the development of robust and generalizable models, which do not suffer from degraded effectiveness when presented with newly-collected, out-of-distribution (OOD) datasets. Contrastive Self-Supervised Learning (SSL) offers a potential solution to labeled data scarcity, as it takes advantage of unlabeled data to increase model effectiveness and robustness. However, the selection of appropriate transformations during the learning process is not a trivial task and often leads to degraded performance or even breaks down the ability of the network to extract meaningful information. In this research, we propose uncovering the optimal augmentations for applying contrastive learning, detecting abnormalities in 1D phonocardiogram (PCG) samples and thus learning a generalized representation of the signal. Specifically, we perform an extensive comparative evaluation of a wide range of audio-based augmentations, evaluate trained classifiers on multiple datasets across different downstream tasks, and finally report on the impact (positive and negative) of each augmentation in model training. We experimentally demonstrate that, depending on its training distribution, the effectiveness of a fully-supervised model can degrade up to 32% when evaluated on unseen data, while SSL models only lose up to 10% or even improve in some cases. We argue and experimentally demonstrate that, contrastive SSL pretraining can assist in providing robust classifiers which can generalize to unseen, OOD data, without relying on time- and labor-intensive annotation processes by medical experts. Furthermore, the proposed extensive evaluation protocol sheds light on the most promising and appropriate augmentations for robust PCG signal processing, by calculating their effect size on model training. Finally, we provide researchers and practitioners with a roadmap towards producing robust models for PCG classification, in addition to an open-source codebase for developing novel approaches.
keywords:
@@
[…]
[inst1]organization=Department of Informatics and Telematics, Harokopio University of Athens,addressline=Omirou 9, city=Tavros, postcode=177 78, state=Athens, country=Greece
[…]
[inst2]organization=Department of Biosciences and Nutrition, Karolinska Institute,addressline=Blickagangen 16, city=Flemingsberg, postcode=141 52, state=Stockholm, country=Sweden
@@ 1 Introduction
Abnormal phonocardiogram (PCG) signals are often an important indicator of heart diseases or deficiencies linked to increased mortality gbd_2017_causes_of_death_collaborators_global_2018 . The detection of such indicators in a non-invasive, accurate, and robust manner could assist towards prevention or timely treatment of such diseases, especially when access to healthcare services is limited.
Automated methods for detecting abnormal samples commonly rely on audio recordings; typically, audio is processed and then used to train classifiers that detect abnormal (i.e., non-healthy) samples and sometimes discriminate between different diseases (e.g., khan2020 ). However, the audio quality of collected signals can be affected by several circumstances. While typical audio-related factors include microphone type or sampling rate, among others, for the case of PCG recordings additional factors can significantly affect the quality of their recording, such as placement of the microphone jaros2023novel ; fontecave2019 , body size, skin, noise giordano2021 ; huisa2023 , etc. The above cases can prove even more challenging, such as when screening an unborn child/fetus, where PCG can only be captured through the mother’s belly kahankova2023 . In addition to the quality of the recording conditions, the training of accurate ML models is heavily dependent on annotated data, the collection of which proves highly time- and cost-consuming. The above is especially important in the biomedical domain, where available samples are scarce to begin with.
As a result, classification models that have been trained on data originating from a single source or on low quality signals fail to generalize to previously unseen data distributions, which do not adhere to the i.i.d. assumption (i.e they are not independently and identically distributed), leading to degraded effectiveness recht2019imagenet ; 10233054 . In this work, we propose training robust classifiers for PCG signal classification by employing contrastive self-supervised learning (SSL) and investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets. However, as the effectiveness of an backbone encoder trained via SSL heavily depends on the adopted augmentation policy Reed_2021_CVPR , we also research the appropriate transformations to be applied on PCG signals and report their effect in model training. Specifically, we attempt to provide insight into the following research questions:
  • 1.
@@
  • 2.
    If so, which augmentations or transformations lead to such representations, proving the most effective, and which actually inhibit training?
  • 3.
@@ 2.1 Deep Learning for Phonocardiogram Classification
The release of publicly available PCG datasets under the scope of the 2016 liu_open_2016 and 2022 reyna_heart_2023 PhysioNet goldberger_physiobank_2000 challenges, has a sparked a line of research in deep learning for PCG classification. The majority of previous works have mainly proposed either 2D or 1D CNN architectures for tackling this particular problem. The main difference lies in whether the proposed models are trained directly on the one dimensional recorded auscultation waves or on spectral images extracted from the audio recordings. For example, multiple recent papers lu2022lightweight ; 10081740 ; 10081866 ; 10081768 propose combining mel-spectrogram images extracted from PCG signals with encoded patient demographic data, in a dual-branch CNN model. The authors find that the combination of the above representations leads to improved accuracy in the heart murmur detection task. Similarly, in another work utilizing frequency domain features, the authors of 10081900 design a hierarchical deep CNN architecture that extracts and combines features from multiple-scales of PCG Mel-Spectrograms. A more recent work, 10122565 , proposes training an ensemble of 15 ResNets He_2016_CVPR with channel-wise attention mechanisms on similar spectrograms, for grading murmur intensity in PCG recordings.
Although few, 1D models also prove promising in the heart-murmur detection domain. In alkhodari_ensemble_2022 the authors employ a transformer-based neural network and train it on 1D wavelet power features extracted from raw PCG signals. On the other hand, 10081735 proposes using a U-Net to predict heart states from the raw PCG signal and combine the predictions with a ResNet model for the final classification. In a different approach, the authors of 10081890 suggest that the combination of a 1D CNN-Bi-LSTM model coupled with attention mechanisms, trained in conjunction with a feature-based classifier leads to robust abnormal PCG signal detection.
In this work, we build upon our preliminary research ballas_listen2yourheart_2022 for the 2022 PhysioNet challenge. In particular, we demonstrate the value of adopting self-supervised contrastive learning for robust PCG classification and also report the top augmentations which should be implemented in such a framework. To our knowledge, this is the first study extensively researching the benefits of each augmentation, and their combinations, for contrastive SSL in 1D PCG signal processing. The aim of this study is to provide future researchers with best-practices regarding PCG signal augmentation, while also propose a structured protocol for the evaluation of developed methods.
@@ 2.2 Self-Supervised Learning in 1D biosignal classification
Self-supervised representation learning (SSL) 9770283 was initially proposed for diminishing the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks misra2020self to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding oord_representation_2019 , the goal of Self-Supervised Contrastive Learning 9462394 ; kumar_contrastive_2022 is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions NEURIPS2021_fa14d4fe . The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which can take advantage of the small amount of data but also yield adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection 9837871 . The authors of mohsenvand_contrastive_2020 propose augmenting multi-channel EEG signals by recombining several channels into additional channels which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying pre-defined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of banville_self-supervised_2019 argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in banville2021uncovering where contrastive predictive coding is implemented on similar SSL-learned features to both predict sleep-stages and detect pathologies in multi-channel EEG signals. Finally, kostas_bendr_2021 proposed BENDR, which combines attention mechanisms vaswani_attention_2017 and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well mehari_self-supervised_2022 , most of which focus on either emotion recognition or pathology detection. For example, in sarkar_self-supervised_2022 the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the art results. In another paper studying emotion recognition from ECG signals, vazquez2022transformer introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal kiyasseh_clocs_2021 , implementing attention mechanisms oh_lead-agnostic_2022 or combining wavelet transformations and random crops of the signal chen_clecg_2021 for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, lan_intra-inter_2022 propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models 9346072 . For example, liu_practical_2023 introduces NeuroPose, a system for 3D hand pose tracking which aims to alleviate the problems imposed due to different sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors lai_contrastive_2022 employ a contrastive loss to pretrain a model on EMG signals for hang-gesture recognition.
@@ 2.3 Model Robustness in 1D biosignal classification
Extracting generalized and robust representations has been one of the most important long-standing goals of machine learning 6472238 . To test and improve the generalizabilty prowess and robustness of deep learning models, several fields have emerged, such as Domain Adaptation wang2018deep , Transfer Learning zhuang2020comprehensive , and Domain Generalization zhou2022domain ; 10233054 ; ballas_multi , which use data from non-i.i.d. distributions or out-of-distribution (OOD) data for model evaluation. These distinct data distributions are referred to as data Domains, as they are most often the outcome of different data generating processes. Here, we reference prior works which fall under the above research fields.
@@
Domain adaptation (DA) methods are the most popular choice when aiming to produce models with improved generalization capabilities in 1D biosignal classification. Similar to Transfer Learning (TL), DA algorithms leverage the representation learning capabilities of models pretrained on diverse source data distributions and fine-tune them on previously unseen but similar target distributions. The end-result is a model which remains robust under the distribution shift between source and target data. In healthcare, DA methods are proposed for a plethora of biosignal classification tasks. In weimann_transfer_2021 , the authors propose pre-training vanilla CNN networks on a large corpus of raw ECG signals and demonstrate their effectiveness when fine-tuned on a smaller dataset for arrhythmia detection. In another more recent work, He et al. he_novel_2023 propose tackling the distribution shift between different arrhythmia detection ECG datasets, by employing three distribution alignment mechanisms. To be more specific, they first extract spatio-temporal features from preprocessed signals, pass them through a graph convolutional neural network and finally align the resulting representations between source and target domains. Many studies have researched the advantages of DA and TL in EEG classification as well wan_review_2021 . To name a few, the authors of 10035017 ; 9154600 and peng_domain_2022 propose aligning source and target distributions via adversarial learning for motor image EEG classification and epilepsy detection respectively, while lan_domain_2018 investigates several state-of-the-art DA algorithms for EEG-based emotion recognition.
@@
Finally, Domain Generalization algorithms aim to tackle the model generalization at its core, as the main goal is to provide generalized models that perform well across unseen datasets. Although limited, there has been a recent interest in biosignal DG research. In the field of PCG classification, the authors of 9298838 propose a ensemble classifier fusion method yields improved results in a DG setup for abnormal vs normal signal classification. To tackle the domain shift and variability present in ECG signals collected from different patients, 9344210 and ma_reducing_2019 propose domain-adversarial model training for ECG and EEG classification, respectively. Most recently, the authors of 10233054 introduce a DG benchmark and a model which leverages representations from multiple layers of the network, inspired from 9898255 , for 12-lead ECG and 64-channel EEG classification.
@@ 3 Materials and Methods
For the self-supervised training of all models, we adapt the SimCLR chen_simple_2020 methodology for 1D PCG representation learning and then evaluate numerous augmentation combinations designed specifically for 1D signal processing tasks.
@@ 3.1 Datasets & Downstream tasks
The PCG signals used in this study originate from a total of 5 distinct public datasets and are available as “wav” files. These include the Ephnogram kazemnejad2021ephnogram , FPCGDB cesarelli_simulation_2012 , Pascal pascal-chsc-2011 , PhysioNet2016 liu_open_2016 , and PhysioNet2022 reyna_heart_2023 databases. Here we describe each dataset separately and provide all necesary details regarding the preprocessing steps followed for the homogenization of all signals into a common format.
@@
The Ephnogram kazemnejad2021ephnogram database contains a total of 69 multi-channel, simultaneous ECG and PCG recordings from 24 healthy adults aged between 23 and 29, acquired during physical activity and rest conditions. Most recordings (62 during exercise) are 30min long, while the duration of the remaining 7 (rest) is about 30 seconds. As specified by the researchers, the dataset includes poor sample signals for noise research purposes. As this dataset was proposed and gathered for multi-modal signal processing purposes and not for pathology detection or classification, it does not include labeled samples. For our study, we use the unlabeled samples of the PCG signals for the SSL pretraining of our models. All PCG recordings were sampled at 8 kHz.
@@
The FPCGDB cesarelli_simulation_2012 or Fetal PCG Database, as its name suggests, is a dataset containing 26 20-minute fetal phonocardiographic signals, gathered from healthy women during their the remaining months of their pregnancy and sampled at 333 Hz. Similarly to the Ephnogram dataset, the collected signals were obtained for the development of fetal PCG signal simulation and fetal heart rate extraction algorithms. Hence, this dataset also does not contain any labeled samples for classification purposes. Furthermore, most recordings are somewhat corrupted by noise, either due to the pregnant human body which contains various noise sources, such as muscular movements or placental blood turbulence, or even due to the surrounding environment.
@@
The Pascal pascal-chsc-2011 Heart Sound Classification challenge was introduced for developing algorithms for the automatic screening of cardiac pathologies. The challenges contains two tasks; one for heart sound segmentation and one for classification. In our work, we use the publicly available labeled data for model training and evaluation which was gathered in both clinical and in-the-wild settings. Notably, the 832 PCG signals are classified as either Normal, Murmur, Extra Heart Sound, Artifact or Extrasystole [^2] Each different category is described in the official challenge description (https://istethoscope.peterjbentley.com/heartchallenge/index.html).. The recording lengths vary from 1 to 30 seconds and are sampled at 195 Hz. As each sample is annotated with a single label, the above task is a multi-class classification problem.
@@
The remaining two datasets were also introduced in the context of public challenges for PCG classification. Similarly to the Pascal challenge, the objective of the PhysioNet2016 liu_open_2016 challenge aims at urging researchers to develop robust algorithms for differentiating normal and abnormal heart sound recordings but instead introduces a binary classification problem. The data included for model training in the dataset contains a total of 3,126 PCG recordings collected from a single precordial location from both healthy subjects and patients suffering from a variety of heart defections, lasting from 5 to around 120 seconds. Once again, the recordings were gathered in both clinical and non-clinical environments, with a sample rate of 2 kHz.
@@
Finally, the PhysioNet2022 challenge reyna_heart_2023 was designed for the development of algorithms which could detect heart murmur in PCG signals. Succedent to the 2016 challenge, the 2022 challenge contains multiple heart sound recordings, collected from multiple auscultation locations. The available data includes 3,163 recordings from 942 patients in total. Each patient has one or more PCG signals, recorded from one of the four heart valves or from a separate prominent auscultation location. Each of the recordings can be classified as ”present”, ”absent” or ”unknown”, meaning that heart murmur presence can be either detected, not detected or can not be determined (e.g due to noise or corrupted audio) respectively, rendering the challenge a multi-class classification problem. The sampling rate of the above signals is 4 kHz.
@@
Each dataset contains signals of varying length, sampling rate and label classes. To accommodate model training in our framework, we follow a similar process as in ballas_listen2yourheart_2022 for converting all signals into a similar format. As a first step, to avoid transient noise artifacts, we initially discard the first and final 2 seconds of each PCG recording. The resulting signal is split into 5 second overlapping windows with a 50% overlap. After resampling all samples to 2 kHz, we assign two types of labels to each window; one that follows the annotations of its original dataset (if it is drawn from a labeled dataset) and one binary label indicating whether it derives from a normal heart sound recording or not [^3] In this study, all recordings initially labeled as unknown are considered abnormal.. For the SSL pretraining stage, we assign pseudolabels to each window, as described in Section 3.2. The assignment of binary labels to each window enables the evaluation of trained models across datasets, which do not necessarily share similar sample classes. For example, a classifier trained to detect normal PCG signals in the Pascal dataset, can be then evaluated without further training on samples from the PhysioNet2022 dataset. The above evaluation process is described in detail in Section 3.4.